Suppose you paid $1,000 for a perpetuity bond that pays $40 a year. Suppose the interest rate drop fall to 1%. What would be the price of the bond if it continues to pay $40 a year?
1
Expert's answer
2019-11-21T09:17:23-0500

If it continues to pay $40 a year, then it's present value is PV = D/r = 40/0.01 = $4,000.